# Sliding DFT-based Signal Recovery for Modulo ADC with 1-bit Folding Information

### v3

@@ II-B Is the 1-bit Folding Information Practical?
One might wonder if generating 1-bit folding information $c[n]$ is practical when implementing a modulo ADC. Early works on modulo ADC hardware have considered recovery methods that utilize the reset count in addition to the self-reset ADC output samples [49, 50]. However, implementing a self-reset ADC with a multi-bit reset count signal may incur a significant increase in power consumption and circuit complexity if the threshold is small [7]. Furthermore, having a multi-bit reset count signal that has the same resolution as the self-reset ADC output samples doubles the storage requirements of the system. Limiting $c[n]$ to 1-bit solves these problems. As shown in Figures 1 and 2, generating $c[n]$ would only require one additional comparator and one adder from the “no c[n]c[n]” modulo ADC setup. The additional circuitry has minimal impact on the overall power consumption, and its penalty is fixed regardless of $b$. For instance, consider a flash ADC architecture for amplitude quantization. An 8-bit modulo ADC with no $c[n]$ would have $2^{8}+2=258$ comparators [^1] The 2 additional comparators come from the comparators of the folding mechanism. whereas an 8-bit modulo ADC with 1-bit folding information $c[n]$ would have 259 comparators. The difference in comparator counts is fixed regardless of the number of bits for amplitude quantization. We also note that the folding information generated in Figure 1 is one bit less than the folding information generated by the modulo ADC hardware proposed in [32]. In [32], the 2-bit signal indicates three states: ‘no crossing’, ‘crossed $\lambda^{\prime}$ threshold’, ‘crossed $\lambda^{\prime}$ threshold’. Finally, we point out that the modulo ADC with 1-bit folding information has already been implemented in hardware [36], but it follows a different architecture from what is shown in Figures 1 and 2.
@@ II-C Signal Recovery Scheme
The proposed recovery algorithm first unfolds $f_{\lambda^{\prime},\mathrm{q}}[n]$ by estimating $z[n]$ and then reconstructs the continuous-time signal from the result. Since the sampling frequency used is above the Nyquist rate, the continuous-time signal can be uniquely identified using an appropriate reconstruction filter. In this subsection, we discuss the unfolding mechanism. An overview of the proposed unfolding procedure is given in Figure 3. The unfolding scheme multiplies a window signal $w[n]$ to small overlapping segments of $f_{\lambda^{\prime},\mathrm{q}}[n]$ and then recovers the modulo residue signal at each segment. The window is slid from left to right until the whole $f_{\lambda^{\prime},\mathrm{q}}[n]$ is processed.
Let $N$ be the number of samples in a short segment of $f_{\lambda^{\prime},\mathrm{q}}[n]$. A length-$N$ tapered cosine window $w[n]$ (a.k.a. Tukey window [51]) can be written as
$w[n]=\begin{cases}\frac{1+\cos\left(\frac{2\pi}{\alpha}\left(\frac{n}{N}-\frac{\alpha}{2}\right)\right)}{2}\;\;\;,\quad 0\leq n<\frac{\alpha N}{2}\\ \qquad\qquad 1\;\quad\qquad,\quad\frac{\alpha N}{2}\leq n<N\left(1-\frac{\alpha}{2}\right)\\ \frac{1+\cos\left(\frac{2\pi}{\alpha}\left(\frac{n}{N}-1+\frac{\alpha}{2}\right)\right)}{2},\quad N\left(1-\frac{\alpha}{2}\right)\leq n<N\end{cases},$
where the parameter $\alpha>0$ is the roll-off factor. We choose $\alpha$ so that $\alpha N\in 2\mathbb{Z}$.
Suppose the $i$-th segment of $f_{\lambda^{\prime},\mathrm{q}}[n]$ is denoted as $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$ and the $i$-th segment of the unfolded signal as $f^{(i)}[n]$. Consecutive segments overlap such that the last $\frac{\alpha N}{2}$ samples of $f_{\lambda^{\prime},\mathrm{q}}^{(i-1)}[n]$ are the first $\frac{\alpha N}{2}$ samples $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$. The window function $w[n]$ is multiplied to the first-order difference of $f_{\lambda^{\prime},\mathrm{q}}[n]$ to get
@@
where $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[-1]=f_{\lambda^{\prime},\mathrm{q}}^{(i-1)}[N-1]\;\forall\;i\in\{2,3,\cdots,I\}$ and we assume that $f_{\lambda^{\prime},\mathrm{q}}^{(1)}[-1]=0$, i.e. signal starts at zero. We now describe the computation of the modulo residue samples in $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$. The principle behind the computation of the modulo residue samples is the Fourier domain separation between the bandlimited signal and the folding instances [10]. The length-$N$ Discrete Fourier Transform (DFT) of $\underline{\tilde{f}_{w}^{(i)}}[n]$ is computed as
@@
Due to the finite bandwidth assumption on $f(t)$, $\hat{F}_{w}(e^{j\frac{2\pi kn}{N}})$ for $\frac{2\pi k}{N}\in(\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}})$ can be written as
@@
where the spectral leakage width, denoted $\delta_{\mathrm{SL}}$, accounts for the frequencies where spectral leakage is still significant. Some energy of the $\underline{\tilde{f}_{w}^{(i)}}[n]$ may leak in the band $(\rho\pi,2\pi-\rho\pi)$ due to the inherent spectral leakage in the DFT computation and cannot be neglected. Spectral leakage vanishes towards zero for high frequencies and cannot be made zero for any signal. However, spectral leakage can be controlled by properly choosing the window length $N$ and roll-off parameter $\alpha$. The signals $\underline{z^{(i)}}[n]$ and $\underline{\epsilon^{(i)}}[n]$ are the first-order difference of the modulo residue samples and quantization noise, respectively, in $\tilde{f}_{w}^{(i)}[n]$. The set $\mathcal{S}_{i}$ contains the indices for which $c[n]=1$ in the $i$-th segment. The set identifies the indices in the $i$-th segment for which the DVG changes its output voltage. The samples $z^{(i)}_{\mathcal{S}_{i}}[n]$ (resp. $z^{(i)}_{\mathcal{S}_{i}^{c}}[n]$) corresponds to the value of the modulo residue for all $n$ such that $c[n]=1$ (resp. $c[n]=0$) in the $i$-th segment. The subscript $\mathrm{OOB}$ in $\hat{F}_{\mathrm{OOB}}$ is used to indicate that this is the set of DFT coefficients outside the desired signal bandwidth plus non-negligible spectral leakage.
We are primarily interested in estimating $f^{(i)}[n]$. Our approach is to estimate $z^{(i)}[n]$ first and then subtract $z^{(i)}[n]$ from $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$. This procedure essentially unfolds the modulo ADC output. Let $K$ be the number of discrete frequencies $k$ such that $\frac{2\pi k}{N}\in(\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}})$. Equation (II-C) can be written in matrix form as
@@
where $\underline{\mathbf{z}_{w}}\in\mathbb{R}^{N\times 1}$, $\underline{\boldsymbol{\epsilon}_{w}}\in\mathbb{R}^{N\times 1}$, and $\hat{\mathbf{F}}_{\mathrm{OOB}}\in\mathbb{C}^{K\times 1}$ are vectorized form of the finite-length signals $w[n]\cdot\underline{z^{(i)}}[n]$, $w[n]\cdot\underline{\epsilon^{(i)}}[n]$, and $\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})$, respectively. The entry of matrix $\mathbf{V}\in\mathbb{C}^{K\times N}$ in the $k^{\prime}$-th row and $n$-th column is $\frac{1}{\sqrt{N}}e^{-j\frac{2\pi nk}{N}}$, where $\frac{2\pi k}{N}$ is the $k^{\prime}$-th discrete frequency in $(\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}})$. Let $\underline{\mathbf{z}_{w,\mathcal{S}_{i}}}\in\mathbb{R}^{|\mathcal{S}_{i}|\times 1}$ and $\underline{\boldsymbol{\epsilon}_{w,\mathcal{S}_{i}}}\in\mathbb{R}^{|\mathcal{S}_{i}|\times 1}$ contain the elements of $\underline{\mathbf{z}_{w}}$ and $\underline{\boldsymbol{\epsilon}_{w}}$ whose indices $n$ are in $\mathcal{S}_{i}$. Then, a pre-estimate of $\underline{\mathbf{z}_{w,\mathcal{S}_{i}}}$, denoted as $\underline{\tilde{\mathbf{z}}_{w,\mathcal{S}_{i}}}$, can be expressed as
$\underline{\tilde{\mathbf{z}}_{w,\mathcal{S}_{i}}}=\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\hat{\mathbf{F}}_{\mathrm{OOB}},$ (7)
where $\mathbf{V}_{\mathcal{S}_{i}}\in\mathbb{C}^{K\times|\mathcal{S}_{i}|}$ is formed by horizontally stacking the columns of $\mathbf{V}$ whose column indices are in $\mathcal{S}_{i}$. We obtain the (windowed) pre-estimate of $\underline{z^{(i)}}[n]$, denoted $\underline{\tilde{z}^{(i)}_{w}}[n]$, by placing the corresponding values of $\underline{\tilde{\mathbf{z}}_{w,\mathcal{S}_{i}}}$ to $\underline{\tilde{z}^{(i)}_{w}}[n]$ for $n\in\mathcal{S}_{i}$. The values of $\underline{\tilde{z}^{(i)}_{w}}[n]$ for $n\notin\mathcal{S}_{i}$ are set to 0.
@@
Note that the $n$-th element of $\underline{\tilde{z}^{(i)}_{w}}[n]$ for $n\in\left\{0,\cdots,\frac{\alpha N}{2}-1\right\}$ has a non-unity scaling of the desired modulo residue signal $\underline{z^{(i)}}[n]$ due to the tapered cosine window $w[n]$. To correct this scaling, we exploit the symmetry of the window function at the edges:
$w[n]+w[n+N\left(1-\frac{\alpha}{2}\right)]=1\quad\forall n\in\{0,1,\cdots,\frac{\alpha N}{2}-1\}.$ (8)
@@
We take the last $\frac{\alpha N}{2}$ samples of $\underline{\tilde{z}^{(i-1)}_{w}}[n]$ and add them to the first $\frac{\alpha N}{2}$ samples of $\underline{\tilde{z}^{(i)}_{w}}[n]$ [^2] Due to the sliding window approach, $\tilde{z}^{(i-1)}_{w}[n]$ is already available when processing the $i$-th segment of $f_{\lambda^{\prime},\mathrm{q}}[n]$. The resulting signal is
@@
The scaling correction is illustrated in Figure 4. Since $z^{(i)}[n]\in 2\lambda^{\prime}\mathbb{Z}$, $\underline{z^{(i)}}[n]\in 2\lambda^{\prime}\mathbb{Z}$. The samples of the first-order difference of modulo residue pre-estimates $\underline{\tilde{z}^{(i)}}[n]$ are rounded to the nearest integer multiple of $2\lambda^{\prime}$ to get the first-order difference of the modulo residue estimates, denoted as $\underline{\hat{z}^{(i)}}[n]$. To undo the first-order difference operation in the modulo residue signal, the formula
@@
is applied. The modulo residue signal estimate $\hat{z}^{(i)}[n]$ is then subtracted from $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$ to unfold the $i$-th segment for $n\in\{0,\cdots,N(1-\frac{\alpha}{2})\}$. The recovery of the unfolded sequence continues until all segments of $f_{\lambda^{\prime},\mathrm{q}}[n]$ have been processed. If $c[n]=0$ for all $n\in\{0,\cdots,N-1\}$ in the $i$-th segment, we can skip the processing of $i$-th segment and proceed to the next one. Finally, a digital lowpass filter with passband region $\left(-\frac{\Omega_{m}}{2}-\delta_{\mathrm{SL}},+\frac{\Omega_{m}}{2}+\delta_{\mathrm{SL}}\right)$ is applied to the result, where $\Omega_{m}=\omega_{m}T_{\mathrm{s}}$. The unfolded signal per segment can be expressed as
@@
$=$ $f^{(i)}[n]+\left\{z^{(i)}[n]-\hat{z}^{(i)}[n]\right\}_{\mathrm{LPF}}+\epsilon^{(i)}_{\mathrm{LPF}}[n].$ (10)
Figure 4: Illustration of the scaling correction for $\underline{\tilde{z}_{w}^{(i)}}[n]$. The top plot shows the folded samples of the modulo ADC for the (i1)(i-1)-th and $i$-th segments, denoted $f_{\lambda^{\prime},\mathrm{q}}^{(i-1)}[n]$ (blue) and $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$ (red), respectively. The yellow and purple lines correspond to modulo thresholds. The bottom plot shows the first-order differences of the (windowed) modulo pre-estimates in the (i1)(i-1)-th and $i$-th segments, denoted $\underline{\tilde{z}_{w}^{(i-1)}}[n]$ (blue) and $\underline{\tilde{z}_{w}^{(i)}}[n]$ (red), respectively. Due to the tapered cosine window, the recovered amplitudes of $\underline{\tilde{z}_{w}^{(i-1)}}[n]$ and $\underline{\tilde{z}_{w}^{(i)}}[n]$ in the overlap region (maroon box) are attenuated. However, adding the last $\alpha N/2$ samples of $\underline{\tilde{z}_{w}^{(i-1)}}[n]$ to the first $\alpha N/2$ samples of $\underline{\tilde{z}_{w}^{(i)}}[n]$ would correct the amplitude of $\underline{\tilde{z}_{w}^{(i)}}[n]$ from $n\in\{0,\cdots,\frac{\alpha N}{2}-1\}$.

### v4

@@ II-B Is the 1-bit Folding Information Practical?
One might wonder whether generating 1-bit folding information $c[n]$ is practical when implementing a modulo ADC. Early works on modulo ADC hardware considered recovery methods that utilize the reset count in addition to the self-reset ADC output samples [42, 44]. However, implementing a self-reset ADC with a multi-bit reset count signal may incur a significant increase in power consumption and circuit complexity, especially when the modulo threshold is small [9]. Furthermore, having a multi-bit reset count signal with the same resolution as the self-reset ADC output samples would double the storage requirements of the system. Limiting $c[n]$ to 1-bit solves these problems.
As shown in Figures 1 and 2, generating $c[n]$ requires only one additional comparator and one adder from the “no c[n]c[n]” modulo ADC setup. The additional circuitry has minimal impact on the overall power consumption, and its penalty is fixed regardless of $b$. For instance, consider a flash ADC architecture for amplitude quantization. An 8-bit modulo ADC with no $c[n]$ would have $2^{8}+2=258$ comparators [^1] The 2 additional comparators come from the comparators of the folding mechanism., whereas an 8-bit modulo ADC with 1-bit folding information $c[n]$ would have 259 comparators. The difference in the comparator counts is fixed regardless of the number of bits for amplitude quantization. We also note that the folding information generated in Figure 1 is one bit less than that generated by the modulo ADC hardware proposed in [30]. In [30], the 2-bit signal indicates three states: ‘no crossing’, ‘crossed $\lambda^{\prime}$ threshold’, ‘crossed $\lambda^{\prime}$ threshold’.
Finally, while the above discussion emphasizes low circuit complexity, it is important to note that our recovery algorithm assumes an ideal timing alignment between $c[n]$ and $f_{\lambda^{\prime},\mathrm{q}}[n]$. In practical implementations, latency, jitter, or other synchronization imperfections can affect the recovery performance. Evaluating the robustness of the STFT-based recovery to such timing offsets is an important topic for future research. Nevertheless, we note that a modulo ADC with 1-bit folding information has already been demonstrated in hardware [46], where both the modulo ADC output and 1-bit folding information are generated using a single clock. This design choice inherently mitigates the timing alignment issues. The specific architecture in [46] differs from that depicted in Figures 1 and 2, but it provides a concrete example showing that generating a synchronized 1-bit folding signal is feasible in practice.
@@ II-C Signal Recovery Scheme
The proposed recovery algorithm first unfolds $f_{\lambda^{\prime},\mathrm{q}}[n]$ by estimating $z[n]$ and then reconstructs the continuous-time signal from the result. Since the sampling frequency used is above the Nyquist rate, the continuous-time signal can be uniquely identified using an appropriate reconstruction filter. In this subsection, we discuss the unfolding mechanism. An overview of the proposed unfolding procedure is shown in Figure 3. The unfolding scheme multiplies a window function $w[n]$ to the small overlapping segments of $f_{\lambda^{\prime},\mathrm{q}}[n]$ and then recovers the modulo residue signal at each segment. The window function is applied sequentially across the signal with a fixed hop size of $N(1-\frac{\alpha}{2})$ until the entire $f_{\lambda^{\prime},\mathrm{q}}[n]$ is processed.
Let $N$ be the number of samples in a short segment of $f_{\lambda^{\prime},\mathrm{q}}[n]$. A length-$N$ tapered cosine window $w[n]$ (also known as Tukey window [34]) can be written as
$w[n]=\begin{cases}\frac{1+\cos\left(\frac{2\pi}{\alpha}\left(\frac{n}{N}-\frac{\alpha}{2}\right)\right)}{2}\;\;\;,\;0\leq n<\frac{\alpha N}{2}\\ \qquad\qquad 1\;\quad\qquad,\;\frac{\alpha N}{2}\leq n<N\left(1-\frac{\alpha}{2}\right)\\ \frac{1+\cos\left(\frac{2\pi}{\alpha}\left(\frac{n}{N}-1+\frac{\alpha}{2}\right)\right)}{2},\;N\left(1-\frac{\alpha}{2}\right)\leq n<N,\end{cases}$ (4)
where the parameter $\alpha>0$ denotes the roll-off factor. We select $\alpha$ such that $\alpha N\in 2\mathbb{Z}$.
Suppose that the $i$-th segment of $f_{\lambda^{\prime},\mathrm{q}}[n]$ is denoted as $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$ and the $i$-th segment of the unfolded signal as $f^{(i)}[n]$. Consecutive segments overlap such that the last $\frac{\alpha N}{2}$ samples of $f_{\lambda^{\prime},\mathrm{q}}^{(i-1)}[n]$ are the first $\frac{\alpha N}{2}$ samples $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$. The window function $w[n]$ is multiplied to the first-order difference of $f_{\lambda^{\prime},\mathrm{q}}[n]$ to obtain
@@
where $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[-1]=f_{\lambda^{\prime},\mathrm{q}}^{(i-1)}[N-1]\;\forall\;i\in\{2,3,\cdots,I\}$ and we assume that $f_{\lambda^{\prime},\mathrm{q}}^{(1)}[-1]=0$, i.e. signal starts at zero. We now describe the computation of the modulo residue samples from $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$. The principle behind the computation of the modulo residue samples is the Fourier domain separation between the bandlimited signal and folding instances [6]. The length-$N$ Discrete Fourier Transform (DFT) of $\underline{\tilde{f}_{w}^{(i)}}[n]$ is computed as
@@
Figure 4: Numerical example illustrating the construction of $\mathbf{V}_{\mathcal{S}_{i}}$ for $N=64$, $\mathrm{OF}=4$, and $\alpha=0.5$. The top-left plot shows the input signal $f(t)$ and the corresponding modulo ADC output $f_{\lambda^{\prime}}[n]$ when $\lambda^{\prime}=0.30$ and no quantization. The folding events are indicated by $c[n]$ in the bottom–left plot. The top–right plot presents the DTFT magnitude spectrum of $f[n]=f(nT_{\mathrm{s}})$ and $w[n]\cdot\underline{f}[n]=w[n]\cdot(f[n]-f[n-1])$. The spectral leakage is visible due to the finite window length. Components below -60 dB are assumed to have negligible impact on the computation. The length-$\delta_{\mathrm{SL}}$ intervals marked in the top-right plot denote the OOB frequency regions whose leakage exceeds -60 dB. These regions correspond to the $K_{\mathrm{SL}}$ discrete frequency indices marked in the DFT spectrum of the (windowed) 1st-order difference of folded signal $w[n]\cdot\underline{f_{\lambda^{\prime}}}[n]$, as shown in the bottom–right plot. The matrix $\mathbf{V}_{\mathcal{S}_{i}}$ is constructed by selecting the rows and columns of a DFT matrix corresponding to the OOB discrete frequency indices and folding indices, respectively.
Due to the finite bandwidth assumption on $f(t)$, $\hat{F}_{w}(e^{j\frac{2\pi k}{N}})$ for $\frac{2\pi k}{N}\in(\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}})$ can be written as
@@
where the spectral leakage width, denoted $\delta_{\mathrm{SL}}$, accounts for the frequencies at which spectral leakage is still significant. Some of the energy of $w[n]\cdot\underline{f^{(i)}}[n]$ may leak in the band $(\rho\pi,2\pi-\rho\pi)$ because of the inherent spectral leakage in the DFT computation and cannot be neglected. Spectral leakage vanishes towards zero at high frequencies and cannot be made zero for any signal. However, the spectral leakage can be controlled by properly choosing the window length $N$ and roll-off parameter $\alpha$. The signals $\underline{z^{(i)}}[n]$ and $\underline{\epsilon^{(i)}}[n]$ are the first-order differences of the modulo residue samples and quantization noise, respectively, in $\tilde{f}_{w}^{(i)}[n]$. The set $\mathcal{S}_{i}$ contains the indices for which $c[n]=1$ in the $i$-th segment. This set identifies the indices in the $i$-th segment for which the DVG changes its output voltage. The signal $z^{(i)}_{\mathcal{S}_{i}}[n]$ (resp. $z^{(i)}_{\mathcal{S}_{i}^{c}}[n]$) corresponds to the modulo residue for all $n$ such that $c[n]=1$ (resp. $c[n]=0$). The subscript $\mathrm{OOB}$ in $\hat{F}_{\mathrm{OOB}}$ indicates that this is the set of DFT coefficients outside the desired signal bandwidth plus non-negligible spectral leakage.
We are primarily interested in estimating $f^{(i)}[n]$. Our approach is to first estimate $z^{(i)}[n]$ and then subtract $z^{(i)}[n]$ from $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$. This procedure essentially unfolds the modulo ADC output. Let $K$ be the number of discrete frequencies $k$ such that $\frac{2\pi k}{N}\in(\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}})$. Equation (II-C) can be written in matrix form as
@@
where $\underline{\mathbf{z}_{w}}\in\mathbb{R}^{N\times 1}$, $\underline{\boldsymbol{\epsilon}_{w}}\in\mathbb{R}^{N\times 1}$, and $\hat{\mathbf{F}}_{\mathrm{OOB}}\in\mathbb{C}^{K\times 1}$ are vectorized form of the finite-length signals $w[n]\cdot\underline{z^{(i)}}[n]$, $w[n]\cdot\underline{\epsilon^{(i)}}[n]$, and $\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})$, respectively. The entry of matrix $\mathbf{V}\in\mathbb{C}^{K\times N}$ in the $k^{\prime}$-th row and $n$-th column is $\frac{1}{\sqrt{N}}e^{-j\frac{2\pi nk}{N}}$, where $\frac{2\pi k}{N}$ is the $k^{\prime}$-th discrete frequency in $(\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}})$. Let $\underline{\mathbf{z}_{w,\mathcal{S}_{i}}}\in\mathbb{R}^{|\mathcal{S}_{i}|\times 1}$ and $\underline{\boldsymbol{\epsilon}_{w,\mathcal{S}_{i}}}\in\mathbb{R}^{|\mathcal{S}_{i}|\times 1}$ denote the subvectors of $\underline{\mathbf{z}_{w}}$ and $\underline{\boldsymbol{\epsilon}_{w}}$, respectively, whose indices $n$ are in $\mathcal{S}_{i}$. We can form the matrix $\mathbf{V}_{\mathcal{S}_{i}}\in\mathbb{C}^{K\times|\mathcal{S}_{i}|}$ by horizontally stacking the columns of $\mathbf{V}$ whose column indices are in $\mathcal{S}_{i}$. A numerical illustration of this construction is provided in Figure 4. The pre-estimate of $\underline{\mathbf{z}_{w,\mathcal{S}_{i}}}$, denoted as $\underline{\tilde{\mathbf{z}}_{w,\mathcal{S}_{i}}}$, is then given by
$\underline{\tilde{\mathbf{z}}_{w,\mathcal{S}_{i}}}=\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\hat{\mathbf{F}}_{\mathrm{OOB}}.$ (9)
Figure 5: Illustration of the scaling correction for $\underline{\tilde{z}_{w}^{(i)}}[n]$. The top plot shows the original signal $f(t)$ and the folded samples of the modulo ADC for the (i1)(i-1)-th and $i$-th segments, denoted as $f_{\lambda^{\prime},\mathrm{q}}^{(i-1)}[n]$ (blue) and $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$ (red), respectively. The dashed constant lines correspond to the modulo thresholds. The bottom plot shows the first-order differences of the (windowed) modulo pre-estimates in the (i1)(i-1)-th and $i$-th segments, denoted $\underline{\tilde{z}_{w}^{(i-1)}}[n]$ (blue) and $\underline{\tilde{z}_{w}^{(i)}}[n]$ (red), respectively. Due to the tapered cosine window, the recovered amplitudes of $\underline{\tilde{z}_{w}^{(i-1)}}[n]$ and $\underline{\tilde{z}_{w}^{(i)}}[n]$ in the overlap region (dash-dot box) are attenuated. However, adding the last $\alpha N/2$ samples of $\underline{\tilde{z}_{w}^{(i-1)}}[n]$ to the first $\alpha N/2$ samples of $\underline{\tilde{z}_{w}^{(i)}}[n]$ corrects the amplitude of $\underline{\tilde{z}_{w}^{(i)}}[n]$.
@@
We obtain the (windowed) pre-estimate of $\underline{z^{(i)}}[n]$, denoted $\underline{\tilde{z}^{(i)}_{w}}[n]$, by placing the corresponding values of $\underline{\tilde{\mathbf{z}}_{w,\mathcal{S}_{i}}}$ to $\underline{\tilde{z}^{(i)}_{w}}[n]$ for $n\in\mathcal{S}_{i}$. The values of $\underline{\tilde{z}^{(i)}_{w}}[n]$ for $n\notin\mathcal{S}_{i}$ are set to 0. Note that the $n$-th element of $\underline{\tilde{z}^{(i)}_{w}}[n]$ for $n\in\left\{0,\cdots,\frac{\alpha N}{2}-1\right\}$ has a non-unity scaling of the desired modulo residue signal $\underline{z^{(i)}}[n]$ due to the tapered cosine window $w[n]$. To correct this scaling, we exploit the symmetry of the window function at the edges for $n\in\{0,1,\cdots,\frac{\alpha N}{2}-1\}$:
$w[n]+w[n+N\left(1-\frac{\alpha}{2}\right)]=1.$ (10)
@@
We take the last $\frac{\alpha N}{2}$ samples of $\underline{\tilde{z}^{(i-1)}_{w}}[n]$ and add them to the first $\frac{\alpha N}{2}$ samples of $\underline{\tilde{z}^{(i)}_{w}}[n]$ [^2] Due to the STFT-based approach, $\tilde{z}^{(i-1)}_{w}[n]$ is already available when processing the $i$-th segment of $f_{\lambda^{\prime},\mathrm{q}}[n]$. The resulting signal is
@@
The scaling correction is illustrated in Figure 5. Since $z^{(i)}[n]\in 2\lambda^{\prime}\mathbb{Z}$, $\underline{z^{(i)}}[n]\in 2\lambda^{\prime}\mathbb{Z}$. The samples of the first-order difference of the modulo residue pre-estimates $\underline{\tilde{z}^{(i)}}[n]$ are rounded to the nearest integer multiple of $2\lambda^{\prime}$ to obtain the first-order difference of the modulo residue estimates, denoted as $\underline{\hat{z}^{(i)}}[n]$. To undo the first-order difference operation in the modulo residue signal, the formula
@@
is applied. The modulo residue signal estimate $\hat{z}^{(i)}[n]$ is then subtracted from $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$ to unfold the $i$-th segment for $n\in\{0,\cdots,N(1-\frac{\alpha}{2})\}$.
The recovery of the unfolded sequence continues until all the segments of $f_{\lambda^{\prime},\mathrm{q}}[n]$ have been processed. If $c[n]=0$ for all $n\in\{0,\cdots,N-1\}$ in the $i$-th segment, we can skip the processing of $i$-th segment and proceed to the next segment. Finally, a digital lowpass filter (LPF) with a passband region $\left(-\frac{\Omega_{m}}{2}-\delta_{\mathrm{SL}},+\frac{\Omega_{m}}{2}+\delta_{\mathrm{SL}}\right)$ is applied to the result, where $\Omega_{m}=\omega_{m}T_{\mathrm{s}}$. The unfolded signal per segment can be expressed as
@@
$=$ $f^{(i)}[n]+\left\{z^{(i)}[n]-\hat{z}^{(i)}[n]\right\}_{\mathrm{LPF}}+\epsilon^{(i)}_{\mathrm{LPF}}[n],$ (12)
where the notation $\{x[n]\}_{\mathrm{LPF}}$ indicates that the signal $x[n]$ is fed to the digital LPF.
II-D Design Considerations for Window Function
Figure 6: Normalized magnitude (in dB) vs. angular frequency of $w[n]\cdot\underline{f}[n]$ for different values of $\alpha$ with (a) $N=64$ and (b) $N=128$. The original signal $f(t)=\sum_{i=0}^{150,000}\mathrm{sinc}(t-i)$ is sampled at four times the Nyquist rate to get $f[n]$. The magnitude response is computed using a max-hold across the zero-padded DFTs of all length-NN frames.
The design of the window function is crucial for the proposed STFT-based recovery algorithm because the unfolding procedure depends on the effective isolation of the out-of-band modulo residue from the in-band signal components. A rectangular window provides the narrowest mainlobe and thus offers the best frequency resolution, but its large sidelobes cause substantial spectral leakage. This leakage is problematic in our setting because it can mask or distort the relatively small out-of-band residue terms used for unfolding. To balance resolution and leakage suppression, we adopt a tapered cosine window. The tapered cosine window preserves a relatively high frequency resolution while offering much lower sidelobes. Moreover, its roll-off parameter allows us to control the trade-off between mainlobe width and sidelobe attenuation. This makes it well suited for separating the in-band and out-of-band components necessary for accurate recovery.
To illustrate this trade-off, we consider the windowing of an oversampled train of sinc pulses. The impact of the roll-off factor $\alpha$ on the spectral leakage of $w[n]\cdot\underline{f}[n]$ for $N=64$ and $N=128$ is shown in Figures 6a and 6b, respectively. As shown in the plots, there is a pronounced difference between the out-of-band energy produced by a rectangular window ($\alpha=0$) and that of a tapered cosine window with $\alpha=0.25$. Increasing $\alpha$ further results in a faster sidelobe decay and improved suppression of out-of-band components.
The choice of the window length $N$ and roll-off parameter $\alpha$ also introduces important design trade-offs. Increasing $N$ improves the frequency resolution but lengthens the STFT frame, thereby increasing the computational complexity of the algorithm. Meanwhile, the roll-off factor $\alpha$ governs how fast the window tapers: larger $\alpha$ yields better sidelobe suppression and reduced leakage, but at the cost of a wider mainlobe and lower frequency resolution; smaller $\alpha$ behaves more like a rectangular window, improving resolution but reintroducing leakage. Furthermore, $\alpha$ dictates the amount of overlap between consecutive segments and the hop size of the algorithm. Thus, the proposed algorithm tends to have a slightly higher computational complexity for a larger $\alpha$ because more overlapping segments must be processed. The impact of $N$ and $\alpha$ on the computational complexity of the proposed algorithm is established in Section III-G. In practice, these window parameters must be chosen jointly to achieve sufficient suppression of spectral leakage while maintaining a manageable computational overhead.